how do you set up a linear function for the following word problem?
Nancy received a piggy bank containing $10.00 and adds an additional $1.00 per week. How much money is in her piggy bank after x weeks?

problem 2
if you use 300 gallons of gas in 4.8 hours, then find the rate of change ($gallons/hour)

Answer by stanbon(75887)

Nancy received a piggy bank containing $10.00 and adds an additional $1.00 per week. How much money is in her piggy bank after x weeks?
Value in x weeks = 10 + x
-----
V(x) = 10 + x
=========================
problem 2
if you use 300 gallons of gas in 4.8 hours, then find the rate of change ($gallons/hour)
-----
rate = (# of gallons)/(# of hours) = 300/4.8 = 62.5 gal/hr
===========================================================
Cheers,
Stan H.
====================